Adult tickets for a play cost $20 and child tickets cost $13. If there were 29 peole at a performance and the theater collected $510 from ticket sales, how many children attended the play?
Answer by Earlsdon(6103)

Let A = the number of adults in attendance and C = the number of children.
From the problem description, you can write:
1) A($20) + C($13) = $510 The cost of adult tickets plus the children's tickets totaled $510
2) A + C = 29 The number of people (adults plus children) totaled 29.
Rewrite equation 2) as: A = 29-C and substitute into equation 1) for A to get:
(29-C)(20) + C(13) = 510 Simplify and solve for C
580-20C + 13C = 510
580-7C = 510 Subtract 580 from both sides.
-7C = -70 Divide both sides by -7
C = 10
10 children attended the play.